a backyard garden is shaped like a rectangle with a length that is 30ft longer than the width and a perimeter of 200ft. what is the w, l and area?
Answer by hyh_123(8)


let w =x
l=x+30
perimeter (p)=200
p=2(l*w)=200
l*w=200/2
l*w=100
put values of l and w
l=x+30 and w=x
x+30+x=100
2x+30 =100
2x =70
x =35
w =35
l =35+30=65
area =l*w
=65*30
=2275